# Holographic duality from Howe duality: Chern–Simons gravity as an ensemble of code CFTs

### v1

@@ 2.2.1 Maximal non-anomalous gaugings
[…]
@@
In fact, this condition can be used to define a maximal non-anomalous subgroup, and the Hilbert space of the new theory after gauging can be defined as the subspace inside $\mathcal{H}$ as the one-dimensional space of wavefunctions satisfying (40).
Generalizing the construction of $\Psi_{\mathcal{C}}$ on $\varSigma$ of a higher genus is straightforward
@@ 2.2.2 Ensemble of gaugings
We would like to compare our setup with that of Benini:2022hzx . There the logic was that a gravitational theory should have no global symmetries Banks:2010zn ; Harlow:2018jwu ; Harlow:2018tng ; Belin:2020jxr , hence to formulate the holographic duality, the one-form symmetry in the bulk must be gauged. Namely one should select a particular maximal non-anomalous subgroup of one-form symmetry, and obtain a “gravitational” theory upon gauging. As we mentioned above, the resulting bulk theory is topologically trivial. It therefore yields a factorized wavefunction when $\varSigma$ is not connected, thus solving the factorization puzzle. Each topologically trivial theory obtained after maximal gauging is holographically dual to a particular Narain CFT, namely code CFT defined in terms of ${\mathcal{C}}$. The holographic duality between a given Narain theory and a topologically trivial theory, which can be described by the $k=1$ Chern-Simons theory (10), was first pointed out in Gukov:2004id , and discussed in detail in Aharony:2023zit .
@@
It is interesting to note that the ensemble of code CFTs defined above has another interpretation as the orbifold groupoid – it is the ensemble of all possible orbifolds of a given Narain theory. The theory is specified by the choice of $\mathcal{O}\in O(n,n,\mathbb{R})$ from (32) and the symmetry that is “orbifolded” is the shift symmetry $({\mathbb{Z}}_{p}\times{\mathbb{Z}}_{p})^{n}\subset U(1)^{n}\times U(1)^{n}$. Orbifolding different subgroups leads to an ensemble of new Narain theories, while the bulk Chern–Simons theory (10), which is the ${\mathbb{Z}}_{p}^{n}$ gauge theory, is the SymTFT of this global symmetry. The resulting construction is the orbifold groupoid of Gaiotto:2020iye .
@@ 2.2.3 Chern–Simons gravity
[…]
@@ 2.3.1 Brief review of codes
$(c,c^{\prime})=\alpha\cdot\beta^{\prime}+\alpha^{\prime}\cdot\beta\ \text{mod }p,$ (44)
@@
A code is self-dual if it agrees with its dual $\mathcal{C}^{\perp}$, defined as the set of all $c^{\prime}\in{\mathscr{D}}$ such that $(c^{\prime},c)=0$ for all $c\in\mathcal{C}$. So we see that the condition for a code to be even is the same as for the Wilson lines with the charges belonging to the subgroup ${\mathcal{C}}\subset{\mathscr{D}}$ to have trivial self-braiding, and that a code is self-dual if it corresponds to a maximal non-anomalous subgroup. The number of even self-dual codes of length $n$ over ${\mathbb{F}}_{p}\times{\mathbb{F}}_{p}$ is Aharony:2023zit
@@
In coding literature, one also defines the complete weight enumerator. For genus one, it is a polynomial in $p^{2}$ variables $x_{\alpha,\beta}$ defined by mapping $\psi_{(\alpha,\beta)}\mapsto\prod_{i}x_{\alpha_{i},\beta_{i}}$. The map from full weight-enumerator to complete weight-enumerator loses partial information about the code as it forgets about the order of the components of the vector $c=(\alpha,\beta)$. In a code CFT construction with rigid embedding the code-blocks $\Psi_{c}$ factorize to a product over $i$ with each factor depending only on the pair $(\alpha_{i},\beta_{i})$. Thus the complete weight-enumerator can be used to determine the CFT partition function, with vanishing fugacities, by substituting each $x_{\alpha,\beta}$ with an appropriate lattice sum Angelinos:2022umf ; Yahagi:2022idq . [^6] Analogous statements are more straightforward for the code CFT construction based on the codes with the sign-definite inner product, e.g. conventional binary codes with Hamming weight relevant for chiral CFTs Dolan:1989kf ; Dolan:1994st ; Henriksson:2021qkt . When formulas for ensemble averages are written in terms of complete weight-enumerators, the sum can be reorganized as a sum over equivalence classes of codes, the contribution from each equivalence class $[\mathcal{C}]$ will be weighted by $\frac{1}{|\operatorname{Aut}[\mathcal{C}]|}$ where $\operatorname{Aut}[\mathcal{C}]$ is the automorphism group.
@@
Codewords $c\in{\mathscr{D}}$ are vectors in ${\mathbb{F}}_{p}^{2n}$. One can consider the group $O^{+}(2n,{\mathbb{F}}_{p}):=O(n,n,{\mathbb{F}}_{p})$ acting on ${\mathbb{F}}_{p}^{2n}$, that leaves the inner product (44) invariant. [^7] The are two nonequivalent orthogonal groups $O^{\pm}(2n,{\mathbb{F}}_{p})$ on ${\mathbb{F}}_{p}^{2n}$ that preserve two different quadratic forms, see appendix B for more details. This group naturally maps codes to codes while preserving their property of being even and/or self-dual. It acts transitively on the set of all even self-dual codes, i.e. maps any given code to any other one. The action on ${\mathbb{F}}_{p}^{2n}$ can be lifted to the vector space of full-weight enumerators $\mathcal{X}$,
@@
The group has $|O^{+}(2n,{\mathbb{F}}_{p})|=(p^{n}-1)p^{n(n-1)}(p^{2}-1)(p^{4}-1)\cdots(p^{2n-2}-1)$ elements.
@@ 2.3.2 Code CFTs
Code-based 2d CFTs were originally constructed from binary error correcting codes in Dolan:1989kf ; Dolan:1994st and more recently in Yahagi:2022idq ; Angelinos:2022umf ; Aharony:2023zit for the class of codes considered here. In short, one considers a Narain CFT Narain:1986am constructed from a lattice of Narain type – even self-dual lattice in ${\mathbb{R}}^{n,n}$ defined by the code via Construction A of Leech and Sloane LeechSloane1971 . The path integral of the resulting CFT is specified by the code as follows. Starting from the code full weight enumerator, one promotes formal variables $\psi_{M}$ for $M=(c_{1},\dots,c_{g})$ from (46) to the codeword blocks $\Psi_{M}$ (31), [^9] For non-zero fugacities, $\Psi_{M}$ is given by (188). We also note, the CFT path integral is different from the CFT partition function by a universal fugacity-dependent factor, see Kraus:2006nb and Aharony:2023zit for details. In what follows we do not make a distinction, while always working with a path integral whenever fugacities are non-zero.
@@
At this point one can readily see that the space $\mathcal{X}$ of enumerator polynomials at genus $g$ is isomorphic to ${\mathcal{H}}^{g}$ – the Hilbert space of CS theory on a genus $g$ surface $\varSigma$, which is equivalent to the tensor product of $g$ copies of ${\mathcal{H}}:={\mathcal{H}}^{1}$,
@@
$T|c\rangle=e^{{i\pi\over k}(c,c)}|c\rangle,\quad S|c\rangle={1\over k^{n}}\sum_{c^{\prime}\in{\mathscr{D}}}e^{-{2\pi i\over k}(c,c^{\prime})}|c^{\prime}\rangle\,.$ (55)
@@
$T\,W[a,b]\,T^{\dagger}=e^{-{i\pi\over k}(b,b)}W[a+b,b]\,,\qquad S\,W[a,b]\,S^{\dagger}=e^{{2\pi i\over k}(a,b)}W[-b,a]\,.$ (56)
@@
Finally, the orthogonal group $O^{+}(2n)$, has the following interpretation in terms of Chern-Simons theory. It is a group of global symmetries implemented by surface operators in the bulk. Its action on ${\mathcal{H}}^{g}$ is given by
@@ 2.4 Qudit formulation
We have seen in section 2.1.2, that the Wilson line operators of the Chern-Simons theory on a genus $g$ Riemann surface $\varSigma$ form the Heisenberg–Weyl group. We will see below that the Hilbert space ${\mathcal{H}}^{g}$ of this theory is a finite-dimensional irreducible representation of the Heisenberg–Weyl group. This is in line with the axiomatic approach to quantum mechanics, where the Hilbert space is often defined, or constructed, as the irrep of the Heisenberg–Weyl group. In what follows we show that ${\mathcal{H}}^{g}$ can be understood as the Hilbert space of $p$-dimensional qudits, giving a quantum information theoretic interpretation to the model of holographic duality discussed in this paper.
@@
We note that $\tau=\zeta^{2^{-1}}$ where $22^{-1}=(p+1)/$ 2 is the inverse element to $2$ in ${\mathbb{F}}_{p}$. It is convenient to define phase space variables ${\rm x}=(a,b)$ and ${\rm y}=(a^{\prime},b^{\prime})$, where ${\rm x,y}\in{\mathbb{F}}_{p}^{2N}$, and a symplectic inner product $[{\rm x},{\rm y}]:=a\cdot b^{\prime}-b\cdot a^{\prime}$. Then it is easy to check that the Weyl operators satisfy the following commutation relation defining the Heisenberg–Weyl group
@@
To see that we define $({\rm x},z):={\mathsf{W}}_{\rm x}\,\zeta^{z}$ for $z\in{\mathbb{F}}_{p}$ and readily find
@@
such that $W[a,b]\leftrightarrow X^{\eta a}Z^{b}$. Here by $\eta\,a$ we mean $\eta_{ij}a^{j}_{I}$, where $1\leq i,j\leq 2n$ and $1\leq I\leq g$. In particular, we find that the Wilson line (33) with charges $a_{I}$ and $b_{I}$ and without the phase factor $\tilde{W}[a,b]:=W[a,b]e^{-{2\pi i\over N}(a,b)}$ corresponds to $(-1)^{(a,b)}{\mathsf{W}}_{\eta a,b}$ This expression is not well defined if charges are understood mod $p$. At the same time, as follows from the definition (33), under the modular group it transforms without any additional phases, c.f. (56)
@@
$S_{A}=\left(\begin{array}[]{cc}(A^{-1})^{T}&0\\ 0&A\end{array}\right),\qquad U_{A}\,|b\rangle=({\rm det}(A))^{{p-1\over 2}}\,|A\,b\rangle,$ (76)
where $b\in F_{p}^{N}$ is understood to be a vector of length $N$. Then for any ${\mathbb{F}}_{p}$-valued and symmetric $T$,
@@
While the combinations of $S_{A},S_{T}$ and $S_{J}$ generate the whole $Sp(2N)$, $U_{A},U_{T},U_{J}$ generate a faithful representation acting on ${\mathcal{H}}_{N}$. The phase factor in the definition of $U_{A}$ (76) is necessary for consistency once the phase factor in $U_{J}$ (82) is introduced: $S_{J}^{2}$ should be the same as $S_{A}$ with $A=-1$. Abolishing these phase factors would result in $U_{A},U_{T},U_{J}$ generating a projective representation – a multiple cover of $Sp(2N)$.
@@
We first note that the equations (70) (with $h$ given by (83)) and (69) are compatible, both being phase-free, because $(-1)^{(b,b)}=1$ for any $b\in{\mathbb{F}}_{p}^{2n}$. Next, we discuss how $U_{h}$ is acting on ${\mathcal{H}}^{g}$. We first consider the “$S$” transformation with $a=0,b=-1,c=1,d=0$. The corresponding $Sp(2N,{\mathbb{F}}_{p})$ matrix can be written as follows $S_{\eta}S_{J}$ where $S_{\eta}$ is the transformation $S_{A}$ (76) with $A=\eta\in GL(2,{\mathbb{F}}_{p})$. Using (76) and (82) we readily find action on the Hilbert space ${\mathcal{H}}_{N}$ to be the Fourier transform

### v2

@@ 2.2.1 Maximal non-anomalous gaugings
The Chern–Simons theory described in the previous section is a bit more complicated. It has a 1-form abelian symmetry group $\mathscr{D}=({\mathbb{Z}}_{k}\times{\mathbb{Z}}_{k})^{n}$. The Wilson lines are parametrized by charges $c=({\alpha},{\beta})\in{\mathscr{D}}$. We call the charge even if it satisfies $\alpha\cdot\beta+\beta\cdot\alpha=0\ ({\rm mod}\,\,2k)$. Wilson lines with even charges are bosonic, i.e. have trivial self-braiding. A subgroup of charges ${\mathcal{C}}\subset{\mathscr{D}}$ has to be non-anomalous to be gauged, meaning all lines with $c\in{\mathcal{C}}$ should be bosonic and have trivial mutual braiding Gaiotto:2014kfa . From (34) one can see that mutual braiding of two lines with charges $c$ and $c^{\prime}$ will be trivial if and only if $\alpha\cdot\beta^{\prime}+\beta\cdot\alpha^{\prime}=0\ ({\rm mod}\,\,k)$. [^6] The non-trivial condition arises when considering the braiding between a line on an $\gamma_{a}$ cycle with a line on the conjugate $\gamma_{b}$ cycle, giving the commutator $W[c,0],W[0,c^{\prime}]]$. Then the condition $\alpha\cdot\beta^{\prime}+\beta\cdot\alpha^{\prime}=0\ ({\rm mod}\,\,k)$ follows from (34).
@@
In fact, this condition can be used to define a maximal non-anomalous subgroup, and the Hilbert space of the new theory after gauging can be defined, as a subspace of $\mathcal{H}$, as the one-dimensional space of wavefunctions satisfying (40).
Generalizing the construction of $\Psi_{\mathcal{C}}$ on $\varSigma$ of a higher genus is straightforward:
@@ 2.2.2 Ensemble of gaugings
We would like to compare our setup with that of Benini:2022hzx . There the logic was that a gravitational theory should have no global symmetries Banks:2010zn ; Harlow:2018jwu ; Harlow:2018tng ; Belin:2020jxr , hence to formulate the holographic duality, the one-form symmetry in the bulk must be gauged. Namely one should select a particular maximal non-anomalous subgroup of one-form symmetry, and obtain a “gravitational” theory upon gauging. As we mentioned above, the resulting bulk theory is topologically trivial. It therefore yields a factorized wavefunction when $\varSigma$ is not connected, thus solving the factorization puzzle. Each topologically trivial theory obtained after maximal gauging is holographically dual to a particular Narain CFT, namely code CFT defined in terms of $\mathcal{C}$}. The holographic duality between a given Narain theory and a topologically trivial theory, which can be described by the $k=1$ Chern–Simons theory (10), was first pointed out in Gukov:2004id , and discussed in detail in Aharony:2023zit .
@@
It is interesting to note that the ensemble of code CFTs defined above has another interpretation as the orbifold groupoid – it is the ensemble of all possible orbifolds of a given Narain theory. The theory is specified by the choice of $\mathcal{O}\in O(n,n,\mathbb{R})$ from (32) and the symmetry that is “orbifolded” is the shift symmetry $({\mathbb{Z}}_{p}\times{\mathbb{Z}}_{p})^{n}\subset U(1)^{n}\times U(1)^{n}$. Orbifolding different subgroups leads to an ensemble of Narain theories, while the bulk Chern–Simons theory (10), which is the ${\mathbb{Z}}_{p}^{n}$ gauge theory, is the SymTFT of this global symmetry. The resulting construction is the orbifold groupoid of Gaiotto:2020iye .
@@ 2.2.3 Chern–Simons gravity
$\begin{matrix}\text{CS with global}\\ \text{1-form symmetry}\end{matrix}\quad\overset{\text{maximal gaugings}}{\longrightarrow}\quad\left\{\begin{matrix}Z_{{\mathcal{C}}_{1}}\\ \vdots\\ Z_{{\mathcal{C}}_{\mathcal{N}}}\end{matrix}\right.\quad\overset{\text{average}}{\longrightarrow}\quad\text{ensemble average}=\text{CS gravity}.$ (42)
@@ 2.3.1 Brief review of codes
[…]
@@
A code is self-dual if it is equal to its dual $\mathcal{C}^{\perp}$, defined as the set of all $c^{\prime}\in{\mathscr{D}}$ such that $(c^{\prime},c)=0$ for all $c\in\mathcal{C}$. So we see that the condition for a code to be even is the same as for the Wilson lines with the charges belonging to the subgroup ${\mathcal{C}}\subset{\mathscr{D}}$ to have trivial self-braiding, and that a code is self-dual if it corresponds to a maximal non-anomalous subgroup. The number of even self-dual codes of length $n$ over ${\mathbb{F}}_{p}\times{\mathbb{F}}_{p}$ is Aharony:2023zit
@@
In coding literature, one also defines the complete weight enumerator. For genus one, it is a polynomial in $p^{2}$ variables $x_{\alpha,\beta}$ defined by mapping $\psi_{(\alpha,\beta)}\mapsto\prod_{i}x_{\alpha_{i},\beta_{i}}$. The map from full weight-enumerator to complete weight-enumerator loses partial information about the code as it forgets about the order of the components of the vector $c=(\alpha,\beta)$. In a code CFT construction with rigid embedding the code-blocks $\Psi_{c}$ factorize to a product over $i$ with each factor depending only on the pair $(\alpha_{i},\beta_{i})$. Thus the complete weight-enumerator can be used to determine the CFT partition function, with vanishing fugacities, by substituting each $x_{\alpha,\beta}$ with an appropriate lattice sum Angelinos:2022umf ; Yahagi:2022idq . [^7] Analogous statements are more straightforward for the code CFT construction based on the codes with the sign-definite inner product, e.g. conventional binary codes with Hamming weight relevant for chiral CFTs Dolan:1989kf ; Dolan:1994st ; Henriksson:2021qkt . When formulas for ensemble averages are written in terms of complete weight-enumerators, the sum can be reorganized as a sum over equivalence classes of codes, the contribution from each equivalence class $\mathcal{C}]$ will be weighted by $\frac{1}{|\operatorname{Aut}[\mathcal{C}]|}$ where $\operatorname{Aut}[\mathcal{C}]$ is the automorphism group.
@@
Codewords $c\in{\mathscr{D}}$ are vectors in ${\mathbb{F}}_{p}^{2n}$. One can consider the group $O^{+}(2n,{\mathbb{F}}_{p}):=O(n,n,{\mathbb{F}}_{p})$ acting on ${\mathbb{F}}_{p}^{2n}$, that leaves the inner product (44) invariant. [^8] The are two inequivalent orthogonal groups $O^{\pm}(2n,{\mathbb{F}}_{p})$ on ${\mathbb{F}}_{p}^{2n}$ that preserve two different quadratic forms, see appendix B for more details. This group naturally maps codes to codes while preserving their property of being even and/or self-dual. It acts transitively on the set of all even self-dual codes, i.e. maps any given code to any other one. The action on ${\mathbb{F}}_{p}^{2n}$ can be lifted to the vector space of full-weight enumerators $\mathcal{X}$,
@@
The group has $|O^{+}(2n,{\mathbb{F}}_{p})|=2(p^{n}-1)p^{n(n-1)}(p^{2}-1)(p^{4}-1)\cdots(p^{2n-2}-1)$ elements.
@@ 2.3.2 Code CFTs
Code-based 2d CFTs were originally constructed from binary error correcting codes in Dolan:1989kf ; Dolan:1994st and more recently in Yahagi:2022idq ; Angelinos:2022umf ; Aharony:2023zit for the class of codes considered here. In short, one considers a Narain CFT Narain:1986am constructed from a lattice of Narain type – even self-dual lattice in ${\mathbb{R}}^{n,n}$ defined by the code via Construction A of Leech and Sloane LeechSloane1971 . The path integral of the resulting CFT is specified by the code as follows. Starting from the code full weight enumerator, one promotes the formal variables $\psi_{M}$ for $M=(c_{1},\dots,c_{g})$ from (46) to the codeword blocks $\Psi_{M}$ (31), [^10] For non-zero fugacities, $\Psi_{M}$ is given by (188). We also note, the CFT path integral is different from the CFT partition function by a universal fugacity-dependent factor, see Kraus:2006nb and Aharony:2023zit for details. In what follows we do not make a distinction, while always working with a path integral whenever fugacities are non-zero.
@@
At this point one can readily see that the space $\mathcal{X}$ of enumerator polynomials at genus $g$ is isomorphic to ${\mathcal{H}}^{g}$ – the Hilbert space of CS theory on a genus-$g$ surface $\varSigma$, which is equivalent to the tensor product of $g$ copies of ${\mathcal{H}}:={\mathcal{H}}^{1}$,
@@
$T|c\rangle=e^{{i\pi\over p}(c,c)}|c\rangle,\quad S|c\rangle={1\over k^{n}}\sum_{c^{\prime}\in{\mathscr{D}}}e^{-{2\pi i\over p}(c,c^{\prime})}|c^{\prime}\rangle\,.$ (55)
@@
$T\,W[a,b]\,T^{\dagger}=e^{-{i\pi\over p}(b,b)}W[a+b,b]\,,\qquad S\,W[a,b]\,S^{\dagger}=e^{{2\pi i\over p}(a,b)}W[-b,a]\,.$ (56)
@@
Finally, the orthogonal group $O^{+}(2n)$, has the following interpretation in terms of Chern–Simons theory. It is a group of global symmetries implemented by surface operators in the bulk. Its action on ${\mathcal{H}}^{g}$ is given by
@@ 2.4 Qudit formulation
We have seen in section 2.1.2 that the Wilson line operators of the Chern–Simons theory on a genus-$g$ Riemann surface $\varSigma$ form the Heisenberg–Weyl group. We will see below that the Hilbert space ${\mathcal{H}}^{g}$ of this theory is a finite-dimensional irreducible representation of the Heisenberg–Weyl group. This is in line with the axiomatic approach to quantum mechanics, where the Hilbert space is often defined, or constructed, as the irrep of the Heisenberg–Weyl group. In what follows we show that ${\mathcal{H}}^{g}$ can be understood as the Hilbert space of $p$-dimensional qudits, giving a quantum information theoretic interpretation to the model of holographic duality discussed in this paper.
@@
We note that $\tau=\zeta^{2^{-1}}$ where $2^{-1}=(p+1)/2$ is the inverse element of $2$ in ${\mathbb{F}}_{p}$. It is convenient to define phase space variables ${\rm x}=(a,b)$ and ${\rm y}=(a^{\prime},b^{\prime})$, where ${\rm x,y}\in{\mathbb{F}}_{p}^{2N}$, and a symplectic inner product ${\rm x},{\rm y}]:=a\cdot b^{\prime}-b\cdot a^{\prime}$. Then it is easy to check that the Weyl operators satisfy the following commutation relation defining the Heisenberg–Weyl group
@@
Defining further $({\rm x},z):={\mathsf{W}}_{\rm x}\,\zeta^{z}$ for $z\in{\mathbb{F}}_{p}$, we readily find
@@
such that $W[a,b]\leftrightarrow X^{\eta a}Z^{b}$. Here by $\eta\,a$ we mean $\eta_{ij}a^{j}_{I}$, where $1\leq i,j\leq 2n$ and $1\leq I\leq g$. In particular, we find that the Wilson line (33) with charges $a_{I}$ and $b_{I}$ and without the phase factor $\tilde{W}[a,b]:=W[a,b]e^{-{2\pi i\over p}(a,b)}$ corresponds to $(-1)^{(a,b)}{\mathsf{W}}_{\eta a,b}$ . This expression is not well defined if charges are understood mod $p$. At the same time, as follows from the definition (33), under the modular group it transforms without any additional phases, c.f. (56)
@@
$S_{A}=\left(\begin{array}[]{cc}(A^{-1})^{T}&0\\ 0&A\end{array}\right),\qquad U_{A}\,|b\rangle=\left(\!\tfrac{\det A}{p}\!\right)\,|A\,b\rangle,$ (76)
where $b\in F_{p}^{N}$ is understood to be a vector of length $N$, and $\left(\!\tfrac{\det A}{p}\!\right)$ is the Legendre symbol of the determinant of AA. [^13] For non-zero $x\in{\mathbb{F}}_{p}$, the Legendre symbol $\left(\!\tfrac{x}{p}\!\right)$ is equal to 1 if $x$ is a square, and 1-1 if it is not. If $A$ implements a permutation $\sigma$, we have simply $\left(\!\tfrac{\det A}{p}\!\right)=(\mathrm{sign}\,\sigma)^{\frac{p-1}{2}}$ for odd primes $p$. Then for any ${\mathbb{F}}_{p}$-valued and symmetric $T$,
@@
While the combinations of $S_{A},S_{T}$ and $S_{J}$ generate the whole $Sp(2N)$, $U_{A},U_{T},U_{J}$ generate a faithful representation acting on ${\mathcal{H}}_{N}$. The phase factor in the definition of $U_{A}$ (76) is necessary for consistency once the phase factor in $U_{J}$ (82) is introduced: $S_{J}^{2}$ should be the same as $S_{A}$ with A=1A=-1. Abolishing these phase factors would result in $U_{A},U_{T},U_{J}$ generating a projective representation – a multiple cover of $Sp(2N)$.
@@
We first note that the equations (70) (with $h$ given by (83)) and (69) are compatible, both being phase-free, because $(-1)^{(b,b)}=1$ for any $b\in{\mathbb{F}}_{p}^{2n}$. Next, we discuss how $U_{h}$ is acting on ${\mathcal{H}}^{g}$. We first consider the “SS” transformation with a=0,b=1,c=1,d=0a=0,b=-1,c=1,d=0. The corresponding $Sp(2N,{\mathbb{F}}_{p})$ matrix can be written as follows $S_{\eta}S_{J}$ where $S_{\eta}$ is the transformation $S_{A}$ (76) with $A=\eta\in GL(2,{\mathbb{F}}_{p})$. Using (76) and (82) we readily find action on the Hilbert space ${\mathcal{H}}_{N}$ to be the Fourier transform